OpenStudy (anonymous):

"Ratios of Areas" the ratio of the perimetes of two similar triangles is 3:5. the scale factor is _____ and the ratio of their area is _____.

OpenStudy (anonymous):

if perimeter is 3:5 the scale factor is 3:5 the area factor is (3/5)^2=9/25=1:(2.8) why ?because area is length*length